OpenStudy (anonymous):

Write the slope-intercept equation of the line perpendicular to y = 5/2 x - 2, which passes though the point (0, 2). Answer y = 5x + 2 y = -2/5x + 2 y = 2x + 5 y = 1/2x + 5

OpenStudy (hba):

Well it is easy First of all find the slope of the equation :)

OpenStudy (anonymous):

perpendicular is opposite reciprical

OpenStudy (anonymous):

pourquoi tu n'as pas de cerveau serieux mon dieu..

OpenStudy (anonymous):

so how do i find slope with the y=mx+b right?

OpenStudy (hba):

Yeah

OpenStudy (anonymous):

what is the slope so far?

OpenStudy (anonymous):

5/2?

OpenStudy (hba):

We will find the slope of this equation and use \[\huge\ m_1 \times m_2=-1\]

OpenStudy (hba):

So we can find the slope of the line perp to it.

OpenStudy (anonymous):

now what is the opposite reciprical of 5/2?

OpenStudy (anonymous):

2/5

OpenStudy (hba):

-2/5 *

OpenStudy (anonymous):

oh so now I?

OpenStudy (hba):

Now use, y-y1=m(x-x1)

OpenStudy (anonymous):

yes -2/5 i say opposite because if the slope was positive the opposite would be negative. n so forth

OpenStudy (hba):

Use the formula i provided :)

OpenStudy (hba):

\[m_2=-1/m_1\]

OpenStudy (anonymous):

so id do the same with this formula like id did the other 1 earlier?

OpenStudy (hba):

Yes :)

OpenStudy (anonymous):

with the 2 ys and 2 xs that was confusing to me??

OpenStudy (hba):

Play the plug in game ;)

OpenStudy (hba):

I am off to sleep @blondie16 will take care of the rest i guess if you need help.

OpenStudy (anonymous):

okay thank you for your help @hba

OpenStudy (anonymous):

so ampy id just get y-2=-2/5(x-5/2)??

OpenStudy (hba):

no x1-0

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

y-y1=m(x-x1) then plug in the y value into the y1 and the x value into the x1 and the new slope into m

OpenStudy (anonymous):

so it would be y-2=-2/5(x-0)?

OpenStudy (anonymous):

other way around xD

OpenStudy (anonymous):

-2/5(x-0)=y-2?

OpenStudy (anonymous):

LOL switch the x and y values

OpenStudy (anonymous):

y-0=-2/5(x-2) ??

OpenStudy (anonymous):

yesh then solve for y... but the answer is pretty easy bc there is only 1 answer with the same slope

OpenStudy (anonymous):

thanks